OpenStudy (anonymous):

If f(4x+3)=2x+1, then find f(x).

OpenStudy (anonymous):

Let y = 4x+3

OpenStudy (anonymous):

f(y) = 2x+1

OpenStudy (anonymous):

4x=y-3 x=(y-3)/4 2x+1 = (y-3)/2 +1

OpenStudy (anonymous):

f(x)=(x-3)/2 +1

OpenStudy (anonymous):

Now give me a best response and close the question! :D